# A non-parametric approach for estimating consumer valuation distributions using second price auctions

### v2

@@ 5.1 Data overview
The data set on eBay on online auctions of Xbox game consoles was obtained from the Modeling Online Auctions data repository. More specifically, we focus on a data set which provides information for $93$ online auctions of identical Xbox game consoles where each auction lasts for $7$ days. For each auction, a user’s bid is recorded only if changes the standing price in the auction. For each such bid, the following information is provided: auctionid (unique auction identifier), bid (dollar value of the bid), bidtime (the time, in days, that the bid was placed), bidder (bidder eBay username), bidderrate (internal eBay rating of the bidder), openbid (the reserve price for the auction, set by the seller), and price (the final selling price for the auction). While the standing price values throughout the course of the auction were not directly provided, they can be easily inferred from the successful bid values from the bid column and the reserve price. Also, the bidtime column directly provides the sequence of times at which there is a change in the standing price.
@@ 5.2 Analysis of Xbox data
Using the Xbox 7-day auctions dataset with slight modifications as mentioned in Section 5.1, we now compute the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$. For the estimation of $\lambda$ (see Section 3.2) we need to choose a subset of auctions whose reserve prices are negligible in the given context. We found that the smallest final selling price in all the auctions is \$25 and the median final selling price in all the auctions is \$125. Given this, we chose all auctions with reserve price less than \$1 ($16$ out of $93$) for obtaining the generalized method of moments based estimator of $\lambda$, and also for the computing the final selling price based estimator $\widehat{F}_{SP}$ (see Step I in Section 3.3). Recall that $\widehat{F}_{SP}$ is one of the components used to compute the initial estimator $\widehat{F}_{init}$.
@@
The plots of the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$ of the (unknown) true valuation distribution along with the corresponding $90\%$ HulC confidence regions are provided in Figure 10. Similar to the phenomenon observed in the simulations in Section 4.2, we notice that the HulC confidence region of $\widehat{F}_{MLE}$ is lesser in width than that of $\widehat{F}_{init}$, indicating comparatively smaller variance of $\widehat{F}_{MLE}$. Another interesting observation is that the curves for these two estimates cross exactly once, with $\widehat{F}_{MLE}(x)$ dominated by $\widehat{F}_{init}(x)$ after the crossing point, and vice-versa before the crossing point. This implies that $\widehat{F}_{init}$ stochastically dominates $\widehat{F}_{MLE}$. In other words, the final selling price/first non-reserve price based initial estimator signifies higher Xbox valuations than the MLE estimator based on the entire collection of standing prices throughout the course of the auctions.
@@
Unlike the simulation setting, the true valuation distribution is obviously not known here. However, we still undertake a limited performance evaluation and comparison exercise for the two approaches. As discussed previously in Section 4.1, if the modeling assumptions are largely unviolated (which seems to be the case) one would expect the MLE to do better than the initial estimator. The goal of this limited evaluation is again to understand the amount of improvement, and also to examine the stability of both estimators. For this purpose, we split the entire Xbox dataset into training and test sets. In particular, we consider two choices of splits namely, $1:1$ and $2:1$, for the ratio of auctions in training vs. test data. For each splitting proportion, 100 random splits are performed. For each split, estimates $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are obtained from the training data, along with the estimate $\widehat{\lambda}_{MLE}$ of the Poisson arrival rate. For each test auction and each of two estimates of $F$ (init and MLE), 100 pseudo auctions with the same reserve price are generated. The mean squared error of the corresponding 100 final selling prices is computed around the observed final selling price in the test auction. These relative mean squared errors are then averaged over all the test auctions. The mean squared errors are finally averaged over these 100 replications/splits to obtain an overall MSE for each of the two methods. The MSEs for both $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are normalized by the MSE of the preliminary estimator $\widehat{F}_{FP}$ (based solely on the first non-reserve standing prices). The resulting normalized MSE values are provided in Table 4. In both settings, MLE yields a better estimate of the final selling price compared to the init.
@@
Figure 10: Plot of Initial estimate (solid blue line, based only on finals selling prices and first observed bids) vs. MLE (solid red line, based on entire sequence of standing prices) of F and their corresponding $90\%$ HulC confidence regions (dotted blue and red lines) for the Xbox 7-day auctions dataset.
@@
Table 4: Mean square error (MSE) between data generated based on reserve price on test set and estimated $\widehat{\lambda}_{MLE}$, $\widehat{\lambda}_{init}$, $\widehat{F}_{MLE,\ test}$ and each of $\widehat{F}_{init,\ training}$, $\widehat{F}_{MLE,\ training}$ and $\widehat{F}_{PT,\ training}$ (with number of bidders generated from MLE model) respectively, compared with data generated from $\widehat{F}_{FP,\ training}$, averaged over $100$ replications of the random split with same proportion, and in each replicate the data is generated 100 times based on the same set of reserve price.
@@ 6 Discussion and future research
In this paper we have a developed a non-parametric methodology for estimating the consumer valuation distribution using second price auction data. Unlike the approach in George and Hui (2012), our methodology uses the collection of current selling price values throughout the course of the auctions, and does not require knowledge of the total number of bidders accessing the auction. Extensive simulations demonstrate that, when the modeling assumptions are true, using our approach can lead to significantly better performance than estimators based on just final selling prices and first observed bids. Two additional assumptions (compared to George and Hui (2012)) which preclude multiple bidding and postulate constant rate of arrival of the consumers to the auction are needed. Many real-life second price auctions see only minor departures from these assumptions, which are supported by economic theory. However, if there is evidence of major violation, results from the proposed methodology should be used cautiously. Generalizing our methodology by relaxing one or both of these assumptions is a topic of future research. One possible direction which we plan on exploring is allow two different rates for the bidder arrival process, with a transition between these two rates happening sometime during the auction period.
Another pertinent direction for future research is incorporation of covariates in the model, see for example P. Bajari and A. Hortacsu (2003). The population valuation distribution for the same product might depend on factors such as condition of the item and seller statistics. One possible path for incorporating this in our model is through a semi-parametric approach where a linear combination of the relevant covariates is added as a location parameter for the cdf F. In particular, the population valuation cdf, evaluated at $t$, with $K$ relevant covariate values $x_{1},x_{2},\ldots,x_{K}$ is given by $F\left(t-\sum_{i=1}^{K}\beta_{i}x_{i}\right)$, see e.g., Groeneboom and Hendrickx (2018). The corresponding likelihood maximization will now involve the unknown regression parameters $\{\beta_{i}\}_{i=1}^{K}$ in addition to $\lambda$ and $F$. While the conditional maximization of $\lambda$ and $F$ will involve similar calculations as those developed in this paper, conditional maximization of the regression coefficients will need more careful thought and analysis.
@@
Rohit Patra’s work was partially supported by NSF grant DMS-2210662.

### v3

@@ 5.1 Data overview
The data set on eBay on online auctions of Xbox game consoles was obtained from the Modeling Online Auctions data repository. More specifically, we focus on a data set which provides information for $93$ online auctions of identical Xbox game consoles where each auction lasts for $7$ days. For each auction, a user’s bid is recorded only if it changes the standing price in the auction. For each such bid, the following information is provided: auctionid (unique auction identifier), bid (dollar value of the bid), bidtime (the time, in days, that the bid was placed), bidder (bidder eBay username), bidderrate (internal eBay rating of the bidder), openbid (the starting price for the auction, set by the seller), and price (the final selling price for the auction). While the standing price values throughout the course of the auction were not directly provided, they can be easily inferred from the successful bid values from the bid column and the starting price. Also, the bidtime column directly provides the sequence of times at which there is a change in the standing price.
@@ 5.2 Analysis of Xbox data
Using the Xbox 7-day auctions dataset with slight modifications as mentioned in Section 5.1, we now compute the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$. For the estimation of $\lambda$ (see Section 3.2) we need to choose a subset of auctions whose starting prices are negligible in the given context. We found that the smallest final selling price in all the auctions is \$25 and the median final selling price in all the auctions is \$125. Given this, we chose all auctions with starting price less than \$1 ($16$ out of $93$) for obtaining the generalized method of moments based estimator of $\lambda$, and also for the computing the final selling price based estimator $\widehat{F}_{SP}$ (see Step I in Section 3.3). Recall that $\widehat{F}_{SP}$ is one of the components used to compute the initial estimator $\widehat{F}_{init}$.
@@
The plots of the initial estimate $\widehat{F}_{init}$ and the MLE $\widehat{F}_{MLE}$ of the (unknown) true valuation distribution along with the corresponding $90\%$ HulC confidence regions are provided in Figure 11. Similar to the phenomenon observed in the simulations in Section 4.2, we notice that the HulC confidence region of $\widehat{F}_{MLE}$ is narrower than that of $\widehat{F}_{init}$, indicating comparatively smaller variance of $\widehat{F}_{MLE}$. Another interesting observation is that the curves for these two estimates cross exactly once, with $\widehat{F}_{MLE}(x)$ dominated by $\widehat{F}_{init}(x)$ after the crossing point, and vice-versa before the crossing point. This implies that $\widehat{F}_{init}$ stochastically dominates $\widehat{F}_{MLE}$. In other words, the final selling price/first non-starting standing price based initial estimator signifies higher Xbox valuations than the MLE estimator based on the entire collection of standing prices throughout the course of the auctions.
@@
Unlike the simulation setting, the true valuation distribution is obviously not known here. However, we still undertake a limited performance evaluation and comparison exercise for the two approaches. As discussed previously in Section 4.1, if the modeling assumptions are largely unviolated (which seems to be the case) one would expect the MLE to do better than the initial estimator. The goal of this limited evaluation is again to understand the amount of improvement, and also to examine the stability of both estimators. For this purpose, we split the entire Xbox dataset into training and test sets. In particular, we consider two choices of splits namely, $1:1$ and $2:1$, for the ratio of auctions in training vs. test data. For each splitting proportion, 100 random splits are performed. For each split, estimates $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are obtained from the training data, along with the estimate $\widehat{\lambda}_{MLE}$ of the Poisson arrival rate. For each test auction and each of two estimates of $F$ (init and MLE), 100 pseudo auctions with the same starting price are generated. The mean squared error of the corresponding 100 final selling prices is computed around the observed final selling price in the test auction. These relative mean squared errors are then averaged over all the test auctions. The mean squared errors are finally averaged over these 100 replications/splits to obtain an overall MSE for each of the two methods. The MSEs for both $\widehat{F}_{MLE}$ and $\widehat{F}_{init}$ are normalized by the MSE of the preliminary estimator $\widehat{F}_{FP}$ (based solely on the first non-starting standing prices). The resulting normalized MSE values are provided in Table 5. In both settings, MLE yields a better estimate of the final selling price compared to the init.
@@
Figure 11: Plot of Initial estimate (solid blue line, based only on final selling prices and first observed bids) vs. MLE (solid red line, based on entire sequence of standing prices) of F and their corresponding $90\%$ HulC confidence regions (dotted blue and red lines) for the Xbox 7-day auctions dataset.
@@
Table 5: Mean squared error (MSE) between data generated based on starting price in the test set and estimated $\widehat{\lambda}_{MLE}$, $\widehat{\lambda}_{init}$, $\widehat{F}_{MLE,\ test}$ and each of $\widehat{F}_{init,\ training}$, $\widehat{F}_{MLE,\ training}$ and $\widehat{F}_{PT,\ training}$ (with number of bidders generated from MLE model) respectively, compared with data generated from $\widehat{F}_{FP,\ training}$, averaged over $100$ replications of the random split with same proportion, and in each replicate the data is generated 100 times based on the same set of starting price.
@@ 6 Discussion and future research
In this paper we have developed a semi-parametric methodology for estimating the consumer valuation distribution using second price auction data. Unlike the approach in George and Hui (2012), our methodology uses the collection of current selling price values throughout the course of the auctions, and does not require knowledge of the total number of bidders accessing the auction. Extensive simulations demonstrate that, when the modeling assumptions are true, using our approach can lead to significantly better performance than estimators based on just final selling prices and first observed bids. Two additional assumptions (compared to George and Hui (2012)) which preclude multiple bidding and postulate constant rate of arrival of the consumers to the auction are needed. Many real-life second price auctions see only minor departures from these assumptions, which are supported by economic theory. However, if there is evidence of major violation, results from the proposed methodology should be used cautiously. Generalizing our methodology by relaxing one or both of these assumptions is a topic of future research. One possible direction which we plan on exploring is allow two different rates for the bidder arrival process, with a transition between these two rates happening sometime during the auction period.
Another pertinent direction for future research is the incorporation of covariates in the model; see, for example, P. Bajari and A. Hortacsu (2003). The population valuation distribution for the same product might depend on factors such as condition of the item and seller statistics. One possible path for incorporating this in our model is through a semi-parametric approach where a linear combination of the relevant covariates is added as a location parameter for the cdf $F$. In particular, the population valuation cdf, evaluated at $t$, with $K$ relevant covariate values $x_{1},x_{2},\ldots,x_{K}$ is given by $F\left(t-\sum_{i=1}^{K}\beta_{i}x_{i}\right)$, see e.g., Groeneboom and Hendrickx (2018). The corresponding likelihood maximization will now involve the unknown regression parameters $\{\beta_{i}\}_{i=1}^{K}$ in addition to $\lambda$ and $F$. While the conditional maximization of $\lambda$ and $F$ will involve similar calculations as those developed in this paper, conditional maximization of the regression coefficients will need more careful thought and analysis.
@@
Rohit Patra’s work was partially supported by NSF grant DMS-2210662. Kshitij Khare’s work was partially supported by NSF grant DMS-2410677.